OpenStudy (anonymous):

The pathway 20m by 2m will be cemented. According to the plan the thickness of the concrete will be 0.15m . if the cost for 1 cubic meter of concrete materials is P1000, what will be the total cost for the materials of the pathway?

OpenStudy (nubeer):

vol = 20 * 2 *0.15 vol = 6m^3 cost = 1000 * 6 cost= 6000